Wednesday 13 July 2011

HBA installation progress — 13-Jul-2011

Today was a challenging day. There had been some rough weather during the 'night' (we are still in 24-hour daylight, of course) and one of the tile covers had lifted. Although the full securing mechanism wasn't in place, it is quite clear that we'll need some additional precautions. As a result we spent part of the morning putting in new connection lines on the tiles... especially the north-west edge. This now means that additional staafjes (connection rods) are required; these have been requested from ASTRON.

We also had some deliveries arrive: three lorries in total. This meant new tiles and lids, but also some additional anchors. However, we still do not have enough anchors (owing to breakages), so again more have been requested from ASTRON.

Loading and unloading has also presented some problems and there was some serious damage on one the tiles. In particular the lower black cover on one tile was punctured on the southeast side, leaving a large gash. The polystyrene underneath was also crushed. We have repaired this using spare parts left over from the winter test tiles. This is not just for the polystyrene, but we added a second lower-cover to that tile for additional support. and to cover the gash. However, the additional weight then presented some new problems and we nearly broke the tile during the lift. This was partly due to the weight, but by this stage we were having additional problems due to high winds that had picked up. Around mid morning, just as we were trying to deploy this tile, the wind speed exceeded 10 metres per second, so we abandoned the attempt.

Work then continued on other areas, such as preparing for the RF-container, levelling and securing the framesets and building and moving other framesets.

The RF-container work has gone well and the location has now been prepared and the concrete foundations are now in place. We've also done some preparation work on the cable ducting, including a prototype cable trough for one of the tiles.

As the afternoon wore on, the wind dropped and we could resume work on the tile deployment. Although we had only managed one tile in the morning, we were able to install the repaired tile by 16:00 hrs and then went on to put in a third. We then returned to the site after dinner to deploy a further three tiles, plus some framesets. The situation on the field is now as follows:

So with six new tiles installed (highlighted with bold white labels), the progress has been good. But it has been tough work and we have been very concerned about the weather conditions. Tomorrow promises to be windy too... but this time with added rain.

No comments:

Post a Comment

Note: only a member of this blog may post a comment.